how do i do substitution on this problem 2x+3y=4, 3x+5y=7 ?
Answer by prabhjyot(164)

2x+3y = 4
3x+5y = 7

Using Substitution Method.
Expressing x as a function of y (from the first equation) x = (4 -3y)/2
Substituting this expression into the second equation:
3(4+3y)/2+5y = 7
Simplifying this expression:
1y = 2
or
y = 2
Now, substituting y = 2 back into the x expressed as a function of y:
x = (4-3y)/2 = -1